OpenStudy (anonymous):

hi, i need help to solve this question about express improper function as partial fraction : (3x^3+6x^2+17x+13) / (x+3)(x^2+4)

OpenStudy (kirbykirby):

You can multiply out the denominator. Then you can use long division to divide the numerator into the denominator, and then use partial fractions on the quotient remainder which will be a proper fraction

OpenStudy (anonymous):

i got the answer, but not pretty sure is it correct or wrong. the answer i get, the questions'equation equal to 3+ -29/9 /(x+3) + (2/9x+13/3)/(x^2 +4) where A= -29/9 , B=2/9, C=13/3, is it correct?